(use present worth analysis) a cost analysis is to be made to determine what, if anything, should be done in a situation where there are three "do something" and one "do nothing" alternatives. estimates of the cost and benefits are as follows:

Alternatives

cost

Uniform annual benefit

End-of-life

Salvage

value

Useful life,

In years

Alt-A

$500

$135

$0

5

Alt-B

600

100

250

5

Alt-C

700

100

180

10

Alt-D

0

0

0

0

Use a 10-year analysis period for the mutually exclusive alternatives. At end of five years, Alternatives 1 and 2 may replaced with identical alternatives (with the same cost, benefits, salvage value, and useful life.

  1. Draw the cash flow diagrams (CFD) for Alt-A, Alt-B and Alt-C
  2. If an 8% interest rate is used, calculate the following:

For Alt-A:

Alt-A: Present Worth of Benefits=                                    

Show the calculations here:

PW of Benefits =

Alt-A: Present Worth of Costs =                                 

Show the calculations here:

PW of costs =

                              

For Alt-B:

Alt-A: Present Worth of Benefits=                                    

Show the calculations here:

PW of Benefits =

Alt-B: Present Worth of Costs =                                 

Show the calculations here:

PW of costs =

For Alt-C:

Alt-A: Present Worth of Benefits=                                    

Show the calculations here:

PW of Benefits =

Alt-C: Present Worth of Costs =                                 

Show the calculations here:

PW of costs =

For Alt-D:

Alt-A: Present Worth of Benefits=                                    

Show the calculations here:

PW of Benefits =

Alt-D: Present Worth of Costs =                                

Show the calculations here:

PW of costs =

  1. Show the calculation for Net Present Worth for each alternative
  2. Selection criteria :
  3. Decision: circle your selected Alternative

Alt-A                Alt-B                Alt-C                Alt-D
Answer #1

The present values have been calculated in the attached picture. Please follow the image and refer to the brief summary below.

Alternative A: Present Value of Benefits = 905.86

Present Value of Cost = 840.29

Net Benefit = 65.57 (PV (Benefit) - PV(Cost)

Alternative B: Please note that I have adjusted the salvage value in the benefits. It can be adjusted in the cost as well. However net benefits remain the same irrespective of that which ultimately determines the best choice

Present Value of Benefits = 956.95

Present Value of Cost = 1008.35

Net Benefit = -51.40 (PV (Benefit) - PV(Cost)

Alternative C: Please note that I have adjusted the salvage value in the benefits. It can be adjusted in the cost as well. However net benefits remain the same irrespective of that which ultimately determines the best choice

Present Value of Benefits = 754.38

Present Value of Cost = 700

Net Benefit = -54.38 (PV (Benefit) - PV(Cost)

Alternative D:

Present Value of Benefits = 0

Present Value of Cost = 0

Net Benefit = 0 (PV (Benefit) - PV(Cost)

Conclusion: Alternative A has the maximum net benefits hence is the optmal choice